English

The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the chance that the problem will be solved

Advertisements
Advertisements

Question

The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the chance that the problem will be solved, if they try independently?

Sum
Advertisements

Solution

Let A be the event that X solves the problem B be the event that Y solves the problem.
Since the odds against student X solving the problem are 8: 6
∴ Probability of occurrence of event A is given by.

P(A) = `6/(8 + 6) = 6/14` and

P(A') = 1 – P(A) = `1 - 6/14 = 8/14`

Also, the odds in favour of student Y solving the problem are 14: 16
∴ Probability of occurrence of event B is given by

P(B) = `14/(14 + 16) = 14/30` and

P(B') = 1 – P(B) = `1 - 14/30 = 16/30`

Now A and B are independent events.
∴ A' and B' are independent events.
A' ∩ B' = Event that neither solves the problem
= P(A' ∩ B') = P(A') · P(B')

= `8/14 xx 16/30`

= `32/105`
A ∪ B = the event that the problem is solved
∴ P (problem will be solved) = P(A ∪ B)
= 1 – P(A ∪ B)`
= 1 – P(A' ∩ B')

= `1 - 32/105`

= `73/105`

shaalaa.com
  Is there an error in this question or solution?
Chapter 7: Probability - Miscellaneous Exercise 7 [Page 110]

APPEARS IN

Balbharati Mathematics and Statistics (Commerce) Part 2 [English] Standard 11 Maharashtra State Board
Chapter 7 Probability
Miscellaneous Exercise 7 | Q 14. (a) | Page 110

RELATED QUESTIONS

A speaks truth in 60% of the cases, while B in 90% of the cases. In what percent of cases are they likely to contradict each other in stating the same fact? In the cases of contradiction do you think, the statement of B will carry more weight as he speaks truth in more number of cases than A?


Given two independent events A and B such that P (A) = 0.3, P (B) = 0.6. Find 

  1. P (A and B)
  2. P(A and not B)
  3. P(A or B)
  4. P(neither A nor B)

The odds against a certain event are 5: 2 and odds in favour of another independent event are 6: 5. Find the chance that at least one of the events will happen.


The odds against a husband who is 55 years old living till he is 75 is 8: 5 and it is 4: 3 against his wife who is now 48, living till she is 68. Find the probability that the couple will be alive 20 years hence.


The odds against a husband who is 55 years old living till he is 75 is 8: 5 and it is 4: 3 against his wife who is now 48, living till she is 68. Find the probability that at least one of them will be alive 20 years hence.


Two hundred patients who had either Eye surgery or Throat surgery were asked whether they were satisfied or unsatisfied regarding the result of their surgery

The follwoing table summarizes their response:

Surgery Satisfied Unsatisfied Total
Throat 70 25 95
Eye 90 15 105
Total 160 40 200

If one person from the 200 patients is selected at random, determine the probability that the person was satisfied given that the person had Throat surgery.


Solve the following:

If P(A) = `"P"("A"/"B") = 1/5, "P"("B"/"A") = 1/3` the find `"P"("B'"/"A'")`


10% of the bulbs produced in a factory are of red colour and 2% are red and defective. If one bulb is picked up at random, determine the probability of its being defective if it is red.


A and B are two events such that P(A) = `1/2`, P(B) = `1/3` and P(A ∩ B) = `1/4`. Find: `"P"("A"/"B")`


Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: P1 + P2 – 2P1P2 


Let A and B be two events such that P(A) = `3/8`, P(B) = `5/8` and P(A ∪ B) = `3/4`. Then P(A|B).P(A′|B) is equal to ______.


If A and B are independent events, then A′ and B′ are also independent


If A and B are mutually exclusive events, then they will be independent also.


If A and B are two independent events then P(A and B) = P(A).P(B).


If A and B are independent, then P(exactly one of A, B occurs) = P(A)P(B') + P(B)P(A') 


Let A and B be independent events P(A) = 0.3 and P(B) = 0.4. Find P(A ∩ B)


Events A and Bare such that P(A) = `1/2`, P(B) = `7/12` and `P(barA ∪ barB) = 1/4`. Find whether the events A and B are independent or not.


Given two independent events, if the probability that exactly one of them occurs is `26/49` and the probability that none of them occurs is `15/49`, then the probability of more probable of the two events is ______.


Five fair coins are tossed simultaneously. The probability of the events that at least one head comes up is ______.


Given \[P(E\cap F)=\frac{1}{6}\] and \[P(E)P(F)=\frac{1}{6}\], what is the relation between E and F?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×